# The twin peaks of learning neural networks

### v1

@@ 3.3 Estimating the Mean Dimension through Monte Carlo
This expression is also valid in the case of binary i.i.d. inputs, where one can identify the influence functions $\tau_{i}^{2}$ with the discrete derivatives:
@@ 3.4 Boolean Mean Dimension
The estimating procedure based on binary sequences presented in the last section does not, in general, lead to the true mean dimension for functions taking real valued inputs from more complex distributions, for example the logits of neural networks trained for image classification. It can be used, however, to efficiently calculate a proxy that we will use it for the rest of the work. In order to distinguish this proxy from the true mean dimension, we call the resulting quantity the Boolean Mean Dimension (BMD). We show in the results below that it can be computed analytically in some cases, and that it empirically correlates with other phenomena in neural networks even in cases where no analytical computation is feasible.
@@ 4.3 Analytical determination of the BMD in the RFM
In Fig. 1 we show the plot of the generalization error and the corresponding BMD of the RFM at a fixed $\alpha_{T}$, as a function of $1/\alpha$ for the MSE (left panels) and CE loss (right panels). As shown in [Gerace et al., 2020], for small regularization $\lambda$ the generalization error has a peak approximately at the interpolation threshold (i.e. where the model starts to fit all the training data). Exactly in the correspondence of the generalization error peak the BMD displays its own peak, meaning that the function implemented by the network is more sensitive to perturbation of the inputs.
@@
Note that in the large dataset limit, when $\alpha,\alpha_{T}\to\infty$ with $\alpha_{D}=\mathcal{O}(1)$, a secondary peak for the BMD of the RFM emerges around $\alpha_{D}=1$, i.e. when the number of parameters of the RFM is the same as the number of input features. This peak is caused by the insurgence of singular values in the spectrum of the covariance matrix $\Omega$ and is more accentuated at lower values of the regularization. In the above plots, in the low regularization regime, this peak is overshadowed by the main BMD peak. We analyze this phenomenology in detail in appendix C.
@@ 5.1 Experimental setup
In the following subsections, each panel displays the performance of a large number of different model architectures with varying degree of over-parameterization, trained on different datasets. Except where specified otherwise, all model are initialized with the common Xavier method [Glorot and Bengio, 2010] and use the Adam optimizer [Kingma and Ba, 2014], with batch size $128$. No specific early stopping criterion is implemented. As in other works analysing the double descent, we experiment with different levels of uniformly random label noise during training, which tends to make the double descent peak more pronounced [Nakkiran et al., 2021]. We discuss the effect of label noise below.
@@ 5.2 MD and generalization peaks as a function of overparametrization
In Fig. 2 we show train and test error, and the BMD for an RFM trained with and without label noise on binary MNIST (even vs odd digits) as a function of the hidden layer width. In Fig. 3, we instead consider a two-layer MLP trained on 10-digits MNIST (varying width) and a ResNet-18 trained on CIFAR10 (varying number of channels), the latter with label noise. In the multi-label case, we are defining the BMD of the network as the average per-logit BMD.
Figure 2: Train error (turquoise), test error (violet) and BMD (orange) curves of the random feature model trained on the MNIST dataset with binary labels on 200 train samples with 0% label noise (left) and 20% label noise (right), tested on 5K samples. The resulting plots represent an average (and standard deviations) obtained repeating 50 different times the experiment.
Figure 3: (Left) Train error (turquoise), test error (violet) and BMD (orange) of the two-layer fully-connected network trained on the MNIST dataset with binary labels on 20K train samples with 15% label noise, tested on the 5K samples. The resulting plot represents an average (and standard deviations) obtained repeating 20 different times the experiment. (Right) Train error, test error and BMD of ResNet-18 trained on the CIFAR-10 with 15% label noise in the train set. The resulting plot represents an average (and standard deviations) obtained repeating 5 different times the experiment.
@@
In all the performed experiments, the BMD displays a peak reached in close correspondence with the interpolation threshold, irrespective of the model class, the training set, or the over-parameterization procedure. Around the critical point, the variance of the model is found to be maximized as the function becomes maximally sensitive to the exact input feature values. Notice that, although the BMD measures the input-output sensitivity of the function implemented by the network with random binary inputs, which are distributed differently compared to the training data, this effect of an enhanced sensitivity around the interpolation threshold is still observable. This is explored further in sub-section 5.6.
@@ 5.3 BMD and Adversarial Initialization
In this section, we analyze the BMD of models under adversarial initialization. This technique can be used to artificially hinder the generalization performance of the model [Liu et al., 2020], but here we explore its effect also on the BMD.
We thus train the RFM in two different initialization settings: in the first setting, we use adversarial initialization and pre-train the model on the MNIST dataset with $100\%$ label noise for $1000$ epochs, followed by training on the original dataset for another $800$ epochs (adversarial initialization). In the second setting, we directly train on the same dataset without label noise for $800$ epochs (non-adversarial initialization). The resulting plot, in Fig. 6 (left panel), represents an average over $10$ different realizations of the experiment. In agreement with the increased generalization error, we observe a significant increase in the BMD of the adversarially initialized model.
Figure 6: (Left): BMD (orange line), train error (turquoise line) and test error (violet line) estimated for the RFM in two different initialization settings: In the first setting, we use adversarial initialization and pretrain the model on the MNIST dataset with 100% label noise for 1000 epochs which is followed by the model being trained with 0% label noise on the same dataset for another 800 epochs (adversarial initialization). In the second setting, we train on the same dataset without label noise for 800 epochs (non-adversarial initialization). The resulting plots represent an average of 10 different realizations of the experiment. (Right): BMD (orange line) and counts (turquoise line) estimated for a two-layer fully connected network trained on the MNIST dataset using 40K train samples with 15% label noise and tested on 110 samples. Counts represent the average amount of sign flips of random pixels of a correctly predicted test image that are necessary to fool the model to a wrong class label. The amount is averaged over all the correctly predicted test data samples. The resulting plot represents an average of 100 different realizations of the experiment. We observe that higher values of the BMD correspond to lower robustness of the model and vice versa.
@@ 5.5 Pixel-Wise Contributions to BMD
The MD as expressed in eq (15) is proportional to a sum of contributions $\tau^{2}_{i}$ of single features indexed by $i$. Similar to [Hahn et al., 2022], we plot these contributions in Fig. 7 as a heatmap, where the bright spots indicate features that contribute strongly to the MD. We show four heatmaps, corresponding to different capacities and at different distances from the BMD peak, for two-layer fully connected networks trained on MNIST.
@@
Figure 7: Heatmaps of the pixel contributions ($\tau^{2}_{i}$ for $1\leq i\leq 784)$ estimated on the two-layer fully connected network trained on the MNIST dataset with 10 classes and normalized to lie within $[0,1]$ interval. The heatmaps correspond to the contributions estimated with respect to label 00 for the models of different capacities (hidden dimensions). The resulting plot represents an average over 20 different runs of the experiment.
@@ 5.6 Different Distributions for Estimating BMD
Figure 8: Mean dimensions of the two-layer fully connected network trained on the MNIST, estimated using Monte Carlo (eq. 16) w.r.t. different distributions. The choice of the distribution does not affect the location of the peak. Moreover, distributions, which first two moments coincide (e.g. binary uniform Unif{$,1\}$ 1 } and Gaussian $\mathcal{N}(0,1)$) yield the same MD pattern. The resulting plot represents an average over 30 different runs of the experiment.
@@ 6 Discussion
Our study raises intriguing questions regarding the potential applications of BMD for regularization purposes. Another interesting future direction could be to investigate how comparing the BMDs achieved by a highly parametrized neural network trained on different datasets can help assess the effective dimensionality of the training data and the complexity of the discriminative tasks.
@@ B.1 Free entropy
and $M_{a}$, $Q_{ab}$ are defined in terms of $q_{ab}$, $p_{ab}$, $r_{a}$ in (B.1). Notice that the integrals inside the “entropic” $G_{S}$ and the $G_{SE}$ tersm can be solved analytically by using multivariate Gaussian integrals identities.
@@
We now restrict the discussion to the case of the convex loss functions as the MSE and the CE loss cited in the main text (4.1). In this case, we have to impose the following scalings
@@ B.2.2 Annealed average of the numerator of the BMD
where we have denoted for simplicity by $\bar{\kappa}_{0}$, $\bar{\kappa}_{1}$, $\bar{\kappa}_{2}$, $\bar{\kappa}_{\star}$ the coefficients in equation (B.1) for $\sigma^{\prime}$. Notice that in all the steps that we have done to arrive performing the average over the inputs we have not used anywhere the binary nature of the inputs. It is therefore easy to see that we would have obtained the same result if we had chosen a probability distribution on the inputs with the same first two moments (e.g., a standard normal distribution). Inserting this back into the previous equation, we get

### v2

@@ 3.3 Estimating the Mean Dimension through Monte Carlo
Note that the definition of the MD for a generic input distribution in Eq. (16), entails a resampling procedure that presumes knowledge of the conditional distribution of a pixel given the rest of the pixel values. In the general case, this pixel is to be resampled multiple times from this conditional distribution, to compute the variance of the function under this variation of the input. This conditional distribution, however, is not a known quantity for a real dataset. For this reason, for example, some authors have proposed an “exchange” procedure, where one randomply samples a different pixel value observed in the same dataset [Hahn et al., 2022], however this approximation neglects the within sample correlations.
Expression (16) can be specialized to the case of binary i.i.d. inputs, where one can identify the influence functions $\tau_{i}^{2}$ with the discrete derivatives:
@@ 3.4 Boolean Mean Dimension
In the general case, the underlying input distribution of the training dataset is not known and estimating the MD on this distribution becomes unfeasible. In the present work, we propose employing the estimation procedure presented in the last section, based on binary sequences, as an easily computable proxy of the sensitivity of the neural network function. In order to distinguish this proxy from the mean dimension over the dataset distribution, we call the resulting quantity the Boolean Mean Dimension (BMD). We show in the results below that the BMD can in some cases be computed analytically, and that it is qualitatively related to the generalization phenomenology in neural networks.
@@ 4.3 Analytical determination of the BMD in the RFM
In Fig. 1 we show the plot of the generalization error and the corresponding BMD of the RFM at a fixed $\alpha_{T}$, as a function of $1/\alpha$ for the MSE (left panels) and CE loss (right panels). As shown in [Gerace et al., 2020], for small regularization $\lambda$ the generalization error develops a peak approximately where the model starts to fit all training data. In the case of the MSE loss, this threshold is often called interpolation threshold and it is located at $N=P$. When using the CE loss, this happens when the projected data become linearly separable and the exact location of the threshold strongly depends on the input statistics and features. Exactly in the correspondence of the generalization error peak the BMD displays its own peak, meaning that the function implemented by the network is more sensitive to perturbation of the inputs.
@@
Note that in the large dataset limit, when $\alpha,\alpha_{T}\to\infty$ with $\alpha_{D}=\mathcal{O}(1)$, a secondary peak for the BMD of the RFM emerges around $\alpha_{D}=1$, i.e. when the number of parameters of the RFM is the same as the number of input features. This peak is caused by the insurgence of singular values in the spectrum of the covariance matrix $\Omega$ and is more accentuated at lower values of the regularization. Since modern deep networks operate in a completely different regime from the large dataset limit specified above, we expect this secondary peak not to be visible in realistic settings. For example, in the above plots in the low regularization regime, this peak is overshadowed by the main BMD peak. We analyze this phenomenology in detail in appendix C.
@@ 5.1 Experimental setup
In the following subsections, each panel displays the performance of a large number of different model architectures with varying degree of over-parameterization, trained on different datasets. Except where specified otherwise, all model are initialized with the common Xavier method [Glorot and Bengio, 2010] and use the Adam optimizer [Kingma and Ba, 2014], with batch size $128$ and learning rate $10^{-4}$. No specific early stopping criterion is implemented. As in other works analysing the double descent, we experiment with different levels of uniformly random label noise during training (which is introduced by corrupting a random fraction of labels), which tends to make the double descent peak more pronounced [Nakkiran et al., 2021]. We discuss the effect of label noise below.
@@ 5.2 MD and generalization peaks as a function of overparametrization
In Fig. 2 we show train and test error, and the BMD for an RFM trained with and without label noise on binary MNIST (even vs odd digits) as a function of the hidden layer width. In Fig. 3, we instead consider a two-layer MLP trained on 10-digits MNIST (varying width) and a ResNet-18 trained on CIFAR10 (varying number of channels), both with label noise. In the multi-label case, we are defining the BMD of the network as the average of the BMDs over the classes, where the output of the network is a vector of predicted log-probabilities for each class (i.e. there is a log-softmax activation in the last layer).
Refer to captionRefer to caption
Figure 2: Train error (turquoise), test error (violet) and BMD (orange) curves of the random feature model trained on the MNIST dataset with binary labels on 5K train samples with 0% label noise (left) and 10% label noise (right), tested on 5K samples. The resulting plots represent an average (and standard deviations) obtained repeating 20 different times the experiment.
Refer to captionRefer to caption
Figure 3: (Left) Train error (turquoise), test error (violet) and BMD (orange) of the two-layer fully-connected network trained on the MNIST dataset with 10 labels on 20K train samples with 20% label noise, tested on the 5K samples. The resulting plot represents an average (and standard deviations) obtained repeating 20 different times the experiment. (Right) Train error, test error and BMD of ResNet-18 trained on the CIFAR-10 with 15% label noise in the train set. The resulting plot represents an average (and standard deviations) obtained repeating 5 different times the experiment.
@@
The BMD displays a peak around the point where the number of parameters of the model allows it to reach zero training error, in close correspondence with the generalization error peak. We find this phenomenology to be robust with respect to the model class, the dataset, and the over-parameterization procedure. Notice however, that standard optimizers based on SGD are able to implicitly regularize the trained models and can strongly reduce the peaking behavior, as already observed in the context of double descent. In the presented figures we introduced label-noise, which ensures the presence of over-fitting and is thus able to restore both peaks.
An important observation is that, in order to see this phenomenology, it is not necessary to account for the training input distribution for the evaluation of the MD, which would not be possible in the case of real data. In fact, in the over-fitting regime, it is possible to detect an increased sensitivity of the neural network function for multiple input distributions, including the i.i.d. binary inputs entailed in the BMD evaluation. This is explored further in sub-section 5.6.
@@ 5.3 BMD and Adversarial Initialization
In this section, we analyze the BMD of two-layer fully connected networks under adversarial initialization  [Liu et al., 2020] on the MNIST dataset. This initialization scheme can be used to artificially hinder the generalization performance of the model, forcing it to converge on a bad minimum of the loss. We here aim to show that the initialization has also an effect on the BMD of the model, increasing the sensitivity of the network.
The adversarial initialization protocol works as follows. We train a two-layer fully connected network in two different phases: in the first phase, we push the network towards an adversarial initialization by pretraining the model with $100\%$ label noise for a fixed amount of epochs; in the second phase, we train the model on the original dataset, with no label noise, for $200$ epochs. The resulting plot, in Fig. 6 (left panel), represents an average over $15$ different realizations of the experiment and shows the effect of the length of the pretraining phase on both generalization performance and BMD of the network. In agreement with our analysis, we observe a simultaneous increase of the two metrics when the adversarial initialization phase is longer and the network is driven towards worse generalization.
Refer to captionRefer to caption
Figure 6: (Left): BMD (orange points), and test error (violet points) estimated for a two-layer fully connected network of width $10^{3}$, trained according to the adversarial initialization protocol described in section 5.3 on 20K samples of the MNIST dataset. On the horizontal axis we vary the number of pretraining epochs, and plot the corresponding increase in the generalization error and the BMD of the model after the second learning stage. The points represent an average of 15 different realizations of the experiment. (Right): BMD (orange line) and counts (turquoise line) estimated for a two-layer fully connected network trained on the MNIST dataset using 20K train samples with 20% label noise and tested on 5K samples. Counts represent the average amount of sign flips of random pixels of a correctly predicted test image that are necessary to fool the model to a wrong class label. The amount is averaged over all the correctly predicted test data samples. The resulting plot represents an average of 40 different realizations of the experiment. We observe that higher values of the BMD correspond to lower robustness of the model and vice versa.
@@ 5.5 Pixel-Wise Contributions to BMD
The MD as expressed in eq (15) is proportional to a sum of contributions $\tau^{2}_{i}$ of single features indexed by $i$. Similar to [Hahn et al., 2022], we plot these contributions in Fig. 7 as a heatmap, where the bright spots indicate features that contribute strongly to the MD. We show four heatmaps, corresponding to different capacities and at different distances from the BMD peak, for a two-layer fully connected network trained on MNIST.
@@
Figure 7: Heatmaps of the pixel contributions ($\tau^{2}_{i}$ for $1\leq i\leq 784)$ estimated on the two-layer fully connected network trained on 20K samples of the MNIST dataset with 10 classes, 20% label noise and normalized to lie within $[0,1]$ interval. The (rescaled) participation ratio is defined as $n\times\frac{\sum_{i=1}^{n}\tau_{i}^{2}}{(\sum_{i=1}^{n}\tau_{i})^{2}}$. After the rescaling, a participation ratio of $1$ indicates a uniform distribution of pixel contributions, while a value of $n$ indicates a distribution concentrated over a single pixel. The heatmaps correspond to the contributions estimated with respect to label 00 for the models of different capacities (hidden layer dimensions) and represent only one seed, while the resulting curves on the plot represent an average over 20 different runs of the experiment.
@@ 5.6 Different Distributions for Estimating BMD
Figure 8: Mean dimensions estimated using Monte Carlo (eq. 16) w.r.t. different distributions of the two-layer fully connected network trained on the 20K of MNIST samples with 20% label noise and tested on 5K samples. The MD values are normalized to lie within [0,1] interval. The choice of the distribution does not affect the location of the peak. Moreover, distributions, which first two moments coincide (e.g. binary uniform Unif{$,1\}$ 1 } and Gaussian $\mathcal{N}(0,1)$) yield the same MD pattern. The resulting plot represents an average over 20 different runs of the experiment.
@@ 6 Discussion
Our study raises intriguing questions regarding the potential applications of BMD for regularization purposes. Another interesting future direction could be to investigate how comparing the BMDs achieved by a highly parametrized neural network trained on different datasets can help assess the effective dimensionality of the training data and the complexity of the discriminative tasks. Finally, it could be interesting to extend the study of the BMD in the RFM in the framework of a polynomial teacher model, recently analyzed in [Aguirre-López et al., 2024].
@@ B.1 Free entropy
and $M_{a}$, $Q_{ab}$ are defined in terms of $q_{ab}$, $p_{ab}$, $r_{a}$ in (53). Notice that the integrals inside the “entropic” $G_{S}$ and the $G_{SE}$ tersm can be solved analytically by using multivariate Gaussian integrals identities.
@@
We now restrict the discussion to the case of the convex loss functions as the MSE and the CE loss cited in the main text (23). In this case, we have to impose the following scalings
@@ B.2.2 Annealed average of the numerator of the BMD
where we have denoted for simplicity by $\bar{\kappa}_{0}$, $\bar{\kappa}_{1}$, $\bar{\kappa}_{2}$, $\bar{\kappa}_{\star}$ the coefficients in equation (50) for $\sigma^{\prime}$. Notice that in all the steps that we have done to arrive performing the average over the inputs we have not used anywhere the binary nature of the inputs. It is therefore easy to see that we would have obtained the same result if we had chosen a probability distribution on the inputs with the same first two moments (e.g., a standard normal distribution). Inserting this back into the previous equation, we get
@@ Appendix E BMD and Data Normalization
Appendix F Effect of the label corruption on the train error
In the Fig. 16 we demonstrate the effect of the label corruption of the train data on the train error, which is allowing to explain the difference between the test and train errors for smaller model widths.
Figure 16: Train error curves of the RFM trained on the MNIST dataset with 10 classes on 1000 train samples with 20 % label noise and estimated on 1000 test samples. The plot demonstrates that better performance of the model on the test data rather than on the train data for smaller widths can be explained by a disproportionally high error on the train examples with the corrupted (wrong) labels (green line), which therefore leads to the higher overall train error (blue line), while tested only on the data with non-corrupted labels the train error (yellow line) is comparable to the test error (red line)